The reduction potential for the following non-standard half cell at 298K is______ volts.

2H+ (5.17×10-3M) + 2e- yields H2 (1.08atm)
Answer #1

Sol .

As Reaction :

2H+ ( 5.17 × 10-3 M ) + 2e- <---> H2 ( 1.08 atm )

So, Reaction Quotient = Q

= p(H2) / [H+]2

= 1.08 / (5.17 × 10-3 )2   = 0.04040 × 106

Also , Standard Reduction potential for this half cell

= E° = 0  

So ,

Reduction potential for this half cell

= E = E° - (0.0591 / n) × log(Q)

where n = no. of electrons transferred = 2

So ,

E = 0 - (0.0591/2) × log(0.04040 × 106 )

= - 0.136 V